OpenStudy (anonymous):

I have to write this as a single natural logarithm. 3 ln 3 + 2 ln x

Nnesha (nnesha):

same like log :=)

Nnesha (nnesha):

power rule first :=) it's like backwards :=)

Nnesha (nnesha):

power rule \[\large\rm log_b x^y = y \log_b x\]

OpenStudy (anonymous):

log3 3^2 = 2 log3 3 I am so bad at these

Nnesha (nnesha):

let's deal with log and ln today! so you can i'm really good at it!

OpenStudy (anonymous):

awesome okay

Nnesha (nnesha):

okay so here is an example ln and log same rules apply for both \[\rm 5 \ln y = \ln y^5\] number at front of ln becomes exponent of y

Nnesha (nnesha):